OpenStudy (anonymous):

Yuka buys 6 pounds of apples at $0.50 a pound. She also buys 6 pounds of bananas. Altogether, the fruit costs $12.00. What is the cost per pound for the bananas? A. $2.00 B. $1.00 C. $1.50 D. $0.50

OpenStudy (31356):

ok so Yuka buys 6 pounds of apples at $0.50 a pound. She also buys 6 pounds of bananas. Altogether, the fruit costs $12.00. What is the cost per pound for the bananas? If equations are easier for you, use the information to make an equation. theres 12 pounds total (6+6) so that will be your total (x+x=12) and you know 1 pound of apples is .50 you also know she bought 6 pounds of bananas. so the equation qould be 1.50x6+bx6=12 (b=bananans, and there were 6 pounds total) so solving that equation should get you 1.50 (Ithink you need to divide 9/6 not 6/9) I hope that helps :/ I'm not that good at explaining

OpenStudy (31356):

Medal?

OpenStudy (anonymous):

THANKS

OpenStudy (31356):

Thanks!